Obtain an expression for the fraction of a complete period that a simple harmonic oscillator spends within a small interval delta_x at a position x. What is its physical significance?
1
Expert's answer
2020-03-20T09:40:48-0400

We know that,

The general equation of the simple harmonic motion

"x= A\sin\omega t"

Now, taking the differentiation with respect to t,

"\dfrac{dx}{dt}=A\omega \cos \omega t"

taking the second derivative of the equation

"\dfrac{d^2x}{dt^2}=-A\omega^2\sin \omega t"

when sin will be maximum =1

"\dfrac{d^2x}{dt^2}=-A\omega^2"